News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 21 April 2017 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was executed on Apr 19 to uplink a flight software patch that updates the EPHIN radiation detector channels E150 and E1300 average count thresholds to prevent fixed overflow errors. Additional real-time procedures were executed on Apr 19-20 to dump OBC-A and -B memory as a follow-up to the patch. The dump will be used to update the baseline memory images maintained on the ground.

A Chandra image release was issued on Apr 19 describing observations of the black hole in the large elliptical galaxy NGC 4696. Data from Chandra and other telescopes show evidence for outbursts from the black hole which pump material and energy into its environment about every 5 to 10 million years. For details see: http://chandra.harvard.edu/photo/2017/ngc4696/

The schedule of targets for the next week is shown below and includes an observation of MXB 1659-298, which was accepted as a Target of Opportunity on Mar 17, a follow-up observation of S255IR NIRS, which was accepted as a Target of Opportunity on Dec 9, and an observation of PSR B1259-63, which was accepted as a Director's Discretionary Time Target of Opportunity on Mar 17.



------------------------------------------
Radiation Belts Apr 24 PSRB1259-63 ACIS-I MXB1659-298 ACIS-S J1306+0735 ACIS-S Apr 25 S255IRNIRS3 ACIS-S Radiation Belts Apr 26 SNR0509-68.7 ACIS-S PG1302-102 HRC-I Apr 27 J1306+0735 ACIS-S A1795 ACIS-S A1795 ACIS-S Apr 28 Kepler-444 ACIS-S ArLac HRC-I 2MASSiJ0729087+4008 ACIS-S Radiation Belts Apr 29 Kepler-444 ACIS-S A1795 ACIS-S Abell1914 ACIS-I J1306+0735 ACIS-S Apr 30 GX3+1 ACIS-S/HETG

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports